OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 65 y + x = 7

OpenStudy (anonymous):

@Preetha

OpenStudy (luigi0210):

Use substitution

OpenStudy (anonymous):

(−8, 1) (1, 6) (6, 1) (9, −2)

OpenStudy (luigi0210):

We are here to guide you to your answer, not give you an answer.

OpenStudy (anonymous):

calculate the value of x from second equation and put it into second equation. and try to solve it by own.

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

@gsRttN

OpenStudy (austinl):

\[y^2 + x^2 = 65\] \[y + x = 7\] x = -y + 7 \[y^2 + (-y + 7)^2 = 65\]

OpenStudy (austinl):

solve for y, and then plug it into the other equation.

OpenStudy (anonymous):

a or c

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i am getting two value of each x and y , (-1,8) and other one is (6,1)

OpenStudy (anonymous):

you can directly put value of x or y from second equation into first and can solve it.. OR you can make complete square of first equaiton by adding +2xy and -2xy in first equation, and you will get complete square of x+y and then you can put value of x+y in it. and it will be easy to calculate.